# Modelling Cosmic-Ray Transport: Magnetized versus Unmagnetized Motion in Astrophysical Magnetic Turbulence

### v2

@@ 3.1 Stochastic Differential Equations
We start with describing the motion of field lines and unmagnetized particles by a 3D Langevin equation (Chandrasekhar, 1943; Bian & Li, 2024)
@@
where $\boldsymbol{V}_{s}$ and $\boldsymbol{X}_{s}$ denote velocity and position of the random walker at the generalized time or path parameter $s$. The walker is parametrized by the scattering rate $\theta$ and thermal velocity $v_{\mathrm{rms}}$, and driven by 3D Gaussian white noise $\mathop{}\!\mathrm{d}{\boldsymbol{W}}_{s}$ with the correlation function $\langle\mathop{}\!\mathrm{d}{W}_{i,s}\mathop{}\!\mathrm{d}{W}_{j,s^{\prime}}\rangle=\delta_{ij}\delta(s-s^{\prime})$. The MSD is given by (as shown in Appendix C.1)
@@
with $D_{xx}=\lim_{s\to\infty}{\langle X_{s}^{2}\rangle}/{2s}$. In the following we set $v_{\mathrm{rms}}=1$ for field lines and unmagnetized transport in accordance with the normalization choices $\|\boldsymbol{V}\|=1$ and $\langle\|\boldsymbol{B}\|^{2}\rangle=1$ according to Section 2.
Next, we describe pitch-angle motion of magnetized particles with an Itô stochastic differential equation of the form (Strauss & Effenberger, 2017)
@@
where $s_{t}$ denotes the displacement along a field line at time $t$ and $\mu_{t}\in(-1,1)$, with reflective boundary conditions, denotes the pitch angle cosine. The effective velocity $v_{\mathrm{eff}}$ reflects anisotropies of the pitch-angle-cosine distribution, where $v_{\mathrm{eff}}=1$ indicates an isotropic uniform distribution. The process is driven by uncorrelated Gaussian white noise with $\langle\mathop{}\!\mathrm{d}{W}_{t}\mathop{}\!\mathrm{d}{W}_{t^{\prime}}\rangle=\delta(t-t^{\prime})$ and is characterized by the pitch angle diffusion coefficient
@@
where the MSD of the field line $\langle X_{\mathrm{fl},s}^{2}\rangle$ is given by equation (13) and the MSD of the pitch-angle walker $\langle s_{t}^{2}\rangle$ is given by equation (17). For the purpose of fitting equation (20) to the data, we evaluate the integral numerically with an adaptive Gaussian quadrature rule, however the asymptotic behaviour can be written explicitly as $\langle Y^{2}_{s_{t}}\rangle\underset{t\to 0}{\sim}v_{\mathrm{eff}}^{2}t^{2}/3$ and $\langle Y^{2}_{s_{t}}\rangle\underset{t\to\infty}{\sim}4\lambda_{\mathrm{fl}}\sqrt{v_{\mathrm{eff}}\lambda_{\mu}t}/3$, as shown in Appendix C.2.
Based on these building blocks, we construct the combined stochastic model $\boldsymbol{Z}_{\mathrm{model},t}$. A stochastic walker alternates between magnetized and unmagnetized transport behaviour, where the duration for each segment is sampled from the respective escape-time probability distributions $p(t|t^{\ast}_{\mu})$ and $p(t|t^{\ast}_{\mathrm{scatter}})$. For magnetized transport, we simulate a field line $\boldsymbol{X}_{\mathrm{fl},s}$ according to equation (12) and let the walker diffuse with $s_{t}$ along this field line according to equation (15). This behaviour is parametrized by the field-line and mirror MFPs $\lambda_{\mathrm{fl}}$ and $\lambda_{\mu}$, the effective velocity $v_{\mathrm{eff}}$ and the magnetized mean duration $t_{\mu}^{\ast}$. Note that we simulate a new independent field line for every time the walker enters magnetized behaviour. For unmagnetized transport, we simply simulate random scattering $\boldsymbol{X}_{\mathrm{scatter},t}$ according to equation (12), which is parametrized by the scattering MFP $\lambda_{\mathrm{scatter}}$ and the scattering mean duration $t_{\mathrm{scatter}}^{\ast}$. The procedure is summarized in Algorithm 1.
@@
2:Choose initial magnetization $m\in\{0,1\}$ $\triangleright$ 0: magnetized, 1: unmagnetized
@@ 3.2 Fitting the Model
As noted in Section 2.4, finding connected segments of particle trajectories, which satisfy the desired conditions $\bar{\kappa}\bar{r}_{g}\lessgtr 1$ becomes more difficult with increasing length of these segments. The reason for this is partly physical due to the distribution of field-line curvature, and partly methodical due to our magnetization condition $\bar{\kappa}\bar{r}_{g}\sim 1$ being merely a heuristic. Thus the conditional data $\langle\Delta X_{\tau}^{2}|\bar{\kappa}\bar{r}_{g}\lessgtr 1\rangle$ comes with the caveat, that data points at larger time scales $\tau$ are backed by a much smaller sample size compared to data points at smaller $\tau$. This is illustrated by the relative bin counts in figure 6(b). When fitting our models, we address this issue by truncating the data at a cut-off time scale $\tau_{\mathrm{cutoff}}$, where the choice of $\tau_{\mathrm{cutoff}}$ should be small enough for sufficient statistical significance and large enough to provide a meaningful fit result. For the magnetized case, we choose $\tau_{\mathrm{cutoff}}=1.5\tau_{\mathrm{peak}}$ with $\tau_{\mathrm{peak}}=\mathop{\mathrm{argmax}}_{\tau}\langle\Delta X^{2}_{\tau}|\bar{\kappa}\bar{r}_{g}<1\rangle/2t$ and ensure a minimum relative bin count of $1\%$. Analogously for the unmagnetized case, we choose $\tau_{\mathrm{cutoff}}=1\tau_{\mathrm{peak}}$ with $\tau_{\mathrm{peak}}=\mathop{\mathrm{argmax}}_{\tau}\langle\Delta X^{2}_{\tau}|\bar{\kappa}\bar{r}_{g}>1\rangle/2t$ and ensure a minimum relative bin count of $1\%$ as well. Despite these precautions, a systematic bias likely remains in the data, most notably in the conditional magnetized data, which is dominated at long time scales by a small number of very intense flux tubes. This remaining bias is addressed by the Bayesian optimization of the magnetized mean duration $t^{\ast}_{\mu}$ and MFP $\lambda_{\mu}$ discussed below.
We proceed by fitting the compound subdiffusion model given by equation (20), which is parametrized by $\lambda_{\mathrm{fl}}$$\lambda_{\mu}$ and $v_{\mathrm{eff}}$, to the magnetized data $\langle\Delta X^{2}_{\tau}|\bar{\kappa}\bar{r}_{g}<1\rangle$. Since a clear distinction of the two MFPs requires reliable data at large $\tau$, we argue that large-scale magnetic mirrors mostly operate on the correlation scale of coherent flux tubes and fix $\lambda_{\mathrm{fl}}=\lambda_{\mu}$, thereby reducing the number of free parameters by one. Further, we fit the Langevin model given by equation (13), which is parametrized by $\lambda_{\mathrm{scatter}}$, to the unmagnetized data $\langle\Delta X_{\tau}^{2}|\bar{\kappa}\bar{r}_{g}>1\rangle$. We employ a non-linear least squares method for both cases (Branch et al. 1999). For comparison, the unconditional asymptotic MFP of the unconditional data is estimated as
@@
with the asymptotic diffusion coefficient $D_{\infty}=\lim_{\tau\to\infty}\langle\Delta X_{\tau}^{2}\rangle/2\tau$ and the particle RMS velocity $v_{\mathrm{rms}}^{2}=\frac{3}{2}\lim_{\tau\to 0}\langle\Delta X_{\tau}^{2}\rangle/\tau^{2}$. This choice of $v_{\mathrm{rms}}$ ensures consistency with the Langevin model for the unmagnetized motion. The resulting MFPs are shown in figure 7(a) and discussed in Section 3.3. The effective pitch angle velocities $v_{\mathrm{eff}}$ are shown in figure 7(b), where $v_{\mathrm{eff}}=1$ corresponds to an isotropic uniform pitch angle cosine distribution. We observe $v_{\mathrm{eff}}>1$ for higher particle energies, which is likely due to shorter durations of magnetized segments, thus those particles experience less mirroring, which would isotropize the pitch angles.
@@
Figure 7: (a) Fitted MFPs for conditional magnetized $\lambda_{\mathrm{fl}}$ and unmagnetized transport $\lambda_{\mathrm{scatter}}$, as well as for the unconditional asymptotic case $\lambda_{\mathrm{asymp}}$. As shown in the inset, $\lambda_{\mathrm{asymp}}$ converges to $\lambda_{\mathrm{scatter}}$ for high energies, where our magnetized model is no longer valid. Scalings $\hat{\omega}_{g}^{-1}$ and $\hat{\omega}_{g}^{-1/3}$ are indicated for reference. The field line MFP $\lambda_{\mathrm{fl}}$ is obtained twice: once by naively fitting equation (20) to the biased magnetized MSD (measured), and once by optimizing the unbiased loss function given by equation (23) (optimized). Both values are scaled by a factor $\frac{1}{4}$ in order to simplify comparison with $\lambda_{\mathrm{scatter}}$ and $\lambda_{\mathrm{asymp}}$. (b) Fitted effective velocity of magnetized pitch-angle diffusion. The error bars in both plots for the fitted models are given by 1.96×1.96\ $\times$ the standard error produced by the respective fit routines. The error bars for $\lambda_{\mathrm{asymp}}$ are obtained by taking the mean and 1.96×1.96\ $\times$ the standard deviation over the 8 independent MHD snapshots.
@@
Figure 8: Escape-time probability distributions estimated from the conditional test-particle averages. (a) The magnetized cases exhibit heavier tails than an exponential distribution, indicating that the magnetized motion has some memory. The power-law scaling $t^{-3/2}$ expected for the classical first-passage time distribution of a random walker on a finite line is indicated for reference. (b) The unmagnetized cases closely resemble exponential distributions, indicating a memory-less Markov nature of the unmagnetized motion.
@@
Figure 9: (a) Measured and fitted mean durations $t^{\ast}_{\mu}$ and $t^{\ast}_{\mathrm{scatter}}$. Box-crossing and decoupling time scales are given for reference. The measured magnetized mean duration is much smaller than the expected decoupling time scale due to neglecting the correlation of large-scale flux tubes. On the other hand, the optimized mean magnetized duration is comparable to the decoupling time scale. The error bars of the optimized results are given by 1.96×1.96\ $\times$ the estimated confidence interval of the Bayesian optimization procedure (see Appendix D). (b) Ratio of measured unmagnetized and magnetized mean durations, as a proxy for the volume filling fraction of scattering sites experienced by the test particles. The optimized mean scattering duration is determined by multiplying this measured ratio with the optimized mean magnetized duration.
@@
Next, we estimate the escape-time probability distributions by recording the lengths of observed conditional magnetized and unmagnetized segments into histograms. The densities of the histogram are plotted in figure 8, revealing an exponential shape for the unmagnetized case, which corroborates the view of the unmagnetized motion as a memory-less random walk. On the other hand, the magnetized case roughly resembles a tempered power-law shape, indicating a process with long memory governed by extended flux tubes. When picturing a 1D pitch-angle walker on a field line with finite length, one would expect the classical first-passage time distribution (Krapivsky et al., 2010), which scales as $t^{-3/2}$. This scaling is indicated for reference.
@@
In order to judge whether the results are reasonable, we compare the magnetized mean duration $t_{\mu}^{\ast}$ with the typical box-crossing time scale $\tau_{\mathrm{box}}$ (see figure 6(a)) and the decoupling time scale $\tau_{\mathrm{decouple}}$, which is estimated as the inflection point of the unconditional running diffusion coefficient $\langle\Delta X_{\tau}^{2}\rangle/2\tau$ as
@@
Firstly, since the dominant coherent structures which govern the particle transport are comparable to the fluid correlation scale $l\lesssim L_{u}<L_{\mathrm{box}}$, we expect $t_{\mu}^{\ast}<\tau_{\mathrm{box}}$, i.e., that magnetized motion is clearly separated from the box-crossing time scale . And secondly, we assume that the initially ballistic and transiently subdiffusive behaviour of the unconditional MSD is given by magnetized transport, and asymptotic diffusion is linked to decoupling of particles from coherent field lines, so we expect $t^{\ast}_{\mu}\sim\tau_{\mathrm{decouple}}$. The various time scales are shown in figure 9(a), which confirms the first expectations, but reveals that the measured magnetized mean durations are shorter than expected by an order of magnitude $\tau_{\mathrm{decouple}}/t_{\mu}^{\ast}\sim O(10)$. Additionally, the combined model given by Algorithm 1 does not reproduce the unconditional MSD with these naively measured parameters (not shown).
This severe underestimation is likely caused by the discrepancy between our combined stochastic model, which simulates a new independent stochastic field line for each magnetized segment, and test particles in MHD snapshots, where magnetized transport is governed by a finite number of spatially correlated intense flux tubes. In order to explore the capabilities of our combined model, we search for an optimal effective mean magnetized duration $t^{\ast}_{\mu}$ by means of Bayesian optimization (see Appendix D for details). Specifically, we minimize the loss function
@@
which compares the unconditional MSD of test particles $\langle\Delta X_{\tau}^{2}\rangle$ with the MSD of our combined model $\langle Z^{2}_{\mathrm{model},\tau}|t^{\ast}_{\mu},\lambda_{\mathrm{fl}}\rangle$. We also take the field line MFP $\lambda_{\mathrm{fl}}$ as a tuneable parameter to account for the previously discussed bias of the magnetized data. Further, the effective pitch-angle walker velocity $v_{\mathrm{eff}}$ and scattering MFP $\lambda_{\mathrm{scatter}}$ are taken from the conditional fit results, the pitch angle MFP is fixed as $\lambda_{\mu}=\lambda_{\mathrm{fl}}$, and the unmagnetized mean duration is determined from the fixed measured ratio $t^{\ast}_{\mathrm{scatter}}/t^{\ast}_{\mu}$ (see figure 9(b)), which serves as a proxy for the volume filling fraction of scattering sites.

### v3

@@ 3.1 Stochastic differential equations
We start with describing the motion of field lines and unmagnetised particles by a 3-D Langevin equation (Chandrasekhar, 1943; Bian & Li, 2024)
@@
where $\boldsymbol{V}_{s}$ and $\boldsymbol{X}_{s}$ denote velocity and position of the random walker at the generalised time or path parameter $s$. The walker is parametrised by the scattering rate $\theta$ and thermal velocity $v_{\mathrm{rms}}$, and driven by 3-D Gaussian white noise $\mathop{}\!\mathrm{d}{\boldsymbol{W}}_{s}$ with the correlation function $\langle\mathop{}\!\mathrm{d}{W}_{i,s}\mathop{}\!\mathrm{d}{W}_{j,s^{\prime}}\rangle=\delta_{ij}\delta(s-s^{\prime})$. The MSD is given by (as shown in Appendix C.1)
@@
with $D_{xx}=\lim_{s\to\infty}{\langle X_{s}^{2}\rangle}/{2s}$. In the following, we set $v_{\mathrm{rms}}=1$ for field lines and unmagnetised transport in accordance with the normalisation choices $\|\boldsymbol{V}\|=1$ and $\langle\|\boldsymbol{B}\|^{2}\rangle=1$ according to Section 2.
Next, we describe pitch-angle motion of magnetised particles with an Itô stochastic differential equation of the form (Strauss & Effenberger, 2017)
@@
where $s_{t}$ denotes the displacement along a field line at time $t$ and $\mu_{t}\in(-1,1)$, with reflective boundary conditions, denotes the pitch angle cosine. The effective velocity $v_{\mathrm{eff}}$ reflects anisotropies of the pitch-angle-cosine distribution, where $v_{\mathrm{eff}}=1$ indicates an isotropic uniform distribution. The process is driven by uncorrelated Gaussian white noise with $\langle\mathop{}\!\mathrm{d}{W}_{t}\mathop{}\!\mathrm{d}{W}_{t^{\prime}}\rangle=\delta(t-t^{\prime})$ and is characterised by the pitch angle diffusion coefficient
@@
where the MSD of the field line $\langle X_{\mathrm{fl},s}^{2}\rangle$ is given by equation (13) and the MSD of the pitch-angle walker $\langle s_{t}^{2}\rangle$ is given by equation (17). For the purpose of fitting equation (20) to the data, we evaluate the integral numerically with an adaptive Gaussian quadrature rule; however, the asymptotic behaviour can be written explicitly as $\langle Y^{2}_{s_{t}}\rangle\underset{t\to 0}{\sim}v_{\mathrm{eff}}^{2}t^{2}/3$ and $\langle Y^{2}_{s_{t}}\rangle\underset{t\to\infty}{\sim}4\lambda_{\mathrm{fl}}\sqrt{v_{\mathrm{eff}}\lambda_{\mu}t}/3$, as shown in Appendix C.2.
Based on these building blocks, we construct the combined stochastic model $\boldsymbol{Z}_{\mathrm{model},t}$. A stochastic walker alternates between magnetised and unmagnetised transport behaviour, where the duration for each segment is sampled from the respective escape-time probability distributions $p(t|t^{\ast}_{\mu})$ and $p(t|t^{\ast}_{\mathrm{scatter}})$. For magnetised transport, we simulate a field line $\boldsymbol{X}_{\mathrm{fl},s}$ according to equation (12) and let the walker diffuse with $s_{t}$ along this field line according to equation (15). This behaviour is parametrised by the field-line and mirror MFPs $\lambda_{\mathrm{fl}}$ and $\lambda_{\mu}$, the effective velocity $v_{\mathrm{eff}}$ and the magnetised mean duration $t_{\mu}^{\ast}$. Note that we simulate a new independent field line for every time the walker enters magnetised behaviour. For unmagnetised transport, we simply simulate random scattering $\boldsymbol{X}_{\mathrm{scatter},t}$ according to equation (12), which is parametrised by the scattering MFP $\lambda_{\mathrm{scatter}}$ and the scattering mean duration $t_{\mathrm{scatter}}^{\ast}$. The procedure is summarised in Algorithm 1.
@@
2:Choose initial magnetisation $m\in\{0,1\}$ $\triangleright$ 0: magnetised, 1: unmagnetised
@@ 3.2 Fitting the model
As noted in Section 2.4, finding connected segments of particle trajectories, which satisfy the desired conditions $\bar{\kappa}\bar{r}_{g}\lessgtr 1$ becomes more difficult with increasing length of these segments. The reason for this is partly physical due to the distribution of field-line curvature and partly methodical due to our magnetisation condition $\bar{\kappa}\bar{r}_{g}\sim 1$ being merely a heuristic. Thus the conditional data $\langle\Delta X_{\tau}^{2}|\bar{\kappa}\bar{r}_{g}\lessgtr 1\rangle$ comes with the caveat, that data points at larger time scales $\tau$ are backed by a much smaller sample size compared to data points at smaller $\tau$. This is illustrated by the relative bin counts in figure 6(b). When fitting our models, we address this issue by truncating the data at a cut-off time scale $\tau_{\mathrm{cutoff}}$, where the choice of $\tau_{\mathrm{cutoff}}$ should be small enough for sufficient statistical significance and large enough to provide a meaningful fit result. For the magnetised case, we choose $\tau_{\mathrm{cutoff}}=1.5\tau_{\mathrm{peak}}$ with $\tau_{\mathrm{peak}}=\mathop{\mathrm{argmax}}_{\tau}\langle\Delta X^{2}_{\tau}|\bar{\kappa}\bar{r}_{g}<1\rangle/2t$ and ensure a minimum relative bin count of $1\%$. Analogously for the unmagnetised case, we choose $\tau_{\mathrm{cutoff}}=1\tau_{\mathrm{peak}}$ with $\tau_{\mathrm{peak}}=\mathop{\mathrm{argmax}}_{\tau}\langle\Delta X^{2}_{\tau}|\bar{\kappa}\bar{r}_{g}>1\rangle/2t$ and ensure a minimum relative bin count of $1\%$ as well. Despite these precautions, a systematic bias likely remains in the data, most notably in the conditional magnetised data, which is dominated at long time scales by a small number of very intense flux tubes. This remaining bias is addressed by the Bayesian optimisation of the magnetised mean duration $t^{\ast}_{\mu}$ and MFP $\lambda_{\mu}$ discussed below.
We proceed by fitting the compound subdiffusion model given by equation (20), which is parametrised by $\lambda_{\mathrm{fl}}$$\lambda_{\mu}$ and $v_{\mathrm{eff}}$, to the magnetised data $\langle\Delta X^{2}_{\tau}|\bar{\kappa}\bar{r}_{g}<1\rangle$. Since a clear distinction of the two MFPs requires reliable data at large $\tau$, we argue that large-scale magnetic mirrors mostly operate on the correlation scale of coherent flux tubes and fix $\lambda_{\mathrm{fl}}=\lambda_{\mu}$, thereby reducing the number of free parameters by one. Further, we fit the Langevin model given by equation (13), which is parametrised by $\lambda_{\mathrm{scatter}}$, to the unmagnetised data $\langle\Delta X_{\tau}^{2}|\bar{\kappa}\bar{r}_{g}>1\rangle$. We employ a nonlinear least squares method for both cases (Branch et al. 1999). For comparison, the unconditional asymptotic MFP of the unconditional data is estimated as
@@
with the asymptotic diffusion coefficient $D_{\infty}=\lim_{\tau\to\infty}\langle\Delta X_{\tau}^{2}\rangle/2\tau$ and the particle r.m.s. velocity $v_{\mathrm{rms}}^{2}=\frac{3}{2}\lim_{\tau\to 0}\langle\Delta X_{\tau}^{2}\rangle/\tau^{2}$. This choice of $v_{\mathrm{rms}}$ ensures consistency with the Langevin model for the unmagnetised motion. The resulting MFPs are shown in figure 7(a) and discussed in Section 3.3. The effective pitch angle velocities $v_{\mathrm{eff}}$ are shown in figure 7(b), where $v_{\mathrm{eff}}=1$ corresponds to an isotropic uniform pitch angle cosine distribution. We observe $v_{\mathrm{eff}}>1$ for higher particle energies, which is likely due to shorter durations of magnetised segments, thus those particles experience less mirroring, which would isotropise the pitch angles.
@@
Figure 7: (a) Fitted MFPs for conditional magnetised $\lambda_{\mathrm{fl}}$ and unmagnetised transport $\lambda_{\mathrm{scatter}}$, as well as for the unconditional asymptotic case $\lambda_{\mathrm{asymp}}$. As shown in the inset, $\lambda_{\mathrm{asymp}}$ converges to $\lambda_{\mathrm{scatter}}$ for high energies, where our magnetised model is no longer valid. Scalings $\hat{\omega}_{g}^{-1}$ and $\hat{\omega}_{g}^{-1/3}$ are indicated for reference. The field line MFP $\lambda_{\mathrm{fl}}$ is obtained twice: once by naively fitting equation (20) to the biased magnetised MSD (measured) and once by optimising the unbiased loss function given by equation (23) (optimised). Both values are scaled by a factor $\frac{1}{4}$ to simplify comparison with $\lambda_{\mathrm{scatter}}$ and $\lambda_{\mathrm{asymp}}$. (b) Fitted effective velocity of magnetised pitch-angle diffusion. The error bars in both plots for the fitted models are given by 1.96×1.96\ $\times$ the standard error produced by the respective fit routines. The error bars for $\lambda_{\mathrm{asymp}}$ are obtained by taking the mean and 1.96×1.96\ $\times$ the standard deviation over the eight independent MHD snapshots.
@@
Figure 8: Escape-time probability distributions estimated from the conditional test-particle averages. (a) Magnetised cases exhibit heavier tails than an exponential distribution, indicating that the magnetised motion has some memory. The power-law scaling $t^{-3/2}$ expected for the classical first-passage time distribution of a random walker on a finite line is indicated for reference. (b) Unmagnetised cases closely resemble exponential distributions, indicating a memory-less Markov nature of the unmagnetised motion.
@@
Figure 9: (a) Measured and fitted mean durations $t^{\ast}_{\mu}$ and $t^{\ast}_{\mathrm{scatter}}$. Box-crossing and decoupling time scales are given for reference. The measured magnetised mean duration is much smaller than the expected decoupling time scale due to neglecting the correlation of large-scale flux tubes. However, the optimised mean magnetised duration is comparable to the decoupling time scale. The error bars of the optimised results are given by 1.96×1.96\ $\times$ the estimated confidence interval of the Bayesian optimisation procedure (see Appendix D). (b) Ratio of measured unmagnetised and magnetised mean durations, as a proxy for the volume filling fraction of scattering sites experienced by the test particles. The optimised mean scattering duration is determined by multiplying this measured ratio with the optimised mean magnetised duration.
@@
Next, we estimate the escape-time probability distributions by recording the lengths of observed conditional magnetised and unmagnetised segments into histograms. The densities of the histogram are plotted in figure 8, revealing an exponential shape for the unmagnetised case, which corroborates the view of the unmagnetised motion as a memory-less random walk. On the other hand, the magnetised case roughly resembles a tempered power-law shape, indicating a process with long memory governed by extended flux tubes. When picturing a 1D pitch-angle walker on a field line with finite length, one would expect the classical first-passage time distribution (Krapivsky et al., 2010), which scales as $t^{-3/2}$. This scaling is indicated for reference.
@@
To judge whether the results are reasonable, we compare the magnetised mean duration $t_{\mu}^{\ast}$ with the typical box-crossing time scale $\tau_{\mathrm{box}}$ (see figure 6(a)) and the decoupling time scale $\tau_{\mathrm{decouple}}$, which is estimated as the inflection point of the unconditional running diffusion coefficient $\langle\Delta X_{\tau}^{2}\rangle/2\tau$ as
@@
First, since the dominant coherent structures which govern the particle transport are comparable to the fluid correlation scale $l\lesssim L_{u}<L_{\mathrm{box}}$, we expect $t_{\mu}^{\ast}<\tau_{\mathrm{box}}$, i.e. that magnetised motion is clearly separated from the box-crossing time scale. Second, we assume that the initially ballistic and transiently subdiffusive behaviour of the unconditional MSD is given by magnetised transport, and asymptotic diffusion is linked to decoupling of particles from coherent field lines, so we expect $t^{\ast}_{\mu}\sim\tau_{\mathrm{decouple}}$. The various time scales are shown in figure 9(a), which confirms the first expectations, but reveals that the measured magnetised mean durations are shorter than expected by an order of magnitude $\tau_{\mathrm{decouple}}/t_{\mu}^{\ast}\sim O(10)$. Additionally, the combined model given by Algorithm 1 does not reproduce the unconditional MSD with these naively measured parameters (not shown).
This severe underestimation is likely caused by the discrepancy between our combined stochastic model, which simulates a new independent stochastic field line for each magnetised segment, and test particles in MHD snapshots, where magnetised transport is governed by a finite number of spatially correlated intense flux tubes. To explore the capabilities of our combined model, we search for an optimal effective mean magnetised duration $t^{\ast}_{\mu}$ by means of Bayesian optimisation (see Appendix D for details). Specifically, we minimise the loss function
@@
which compares the unconditional MSD of test particles $\langle\Delta X_{\tau}^{2}\rangle$ with the MSD of our combined model $\langle Z^{2}_{\mathrm{model},\tau}|t^{\ast}_{\mu},\lambda_{\mathrm{fl}}\rangle$. We also take the field line MFP $\lambda_{\mathrm{fl}}$ as a tuneable parameter to account for the previously discussed bias of the magnetised data. Further, the effective pitch-angle walker velocity $v_{\mathrm{eff}}$ and scattering MFP $\lambda_{\mathrm{scatter}}$ are taken from the conditional fit results, the pitch angle MFP is fixed as $\lambda_{\mu}=\lambda_{\mathrm{fl}}$ and the unmagnetised mean duration is determined from the fixed measured ratio $t^{\ast}_{\mathrm{scatter}}/t^{\ast}_{\mu}$ (see figure 9(b)), which serves as a proxy for the volume filling fraction of scattering sites.